could you please show me how to graph -2x+4y+16=0
Answer by Cromlix(4381)

First of all you must transform the equation into the form of a straight line, namely y = mx + c => -2x + 4y + 16 = 0 => 4y = 2x - 16 => Divide through by 4 to reduce the 4y to y => y = 1/2 x - 4 This is the equation.
Next you substitute x values into the equation to generate y values.
x = -2 will give y = -5. x = 0 will give y = -4. x = 2, y = -3. x =4, y = -2.
x = 6, y = -1. x = 8, y = 0. x = 10, y = 1.
You then plot these points on a graph with the 'x' axis horizontal and the y axis vertical.
This will give you a line graph.
Hope this helps.